OpenStudy (anonymous):

Sid lives 3 miles north and 8 miles west of school. What is the direct distance from Sid's house to school?

OpenStudy (anonymous):

 This is our problem, right? Pythagorean theorem will help here.

OpenStudy (anonymous):

11

OpenStudy (anonymous):

\[d^2=3^2+8^2=9+64=75\] So then: \[d=\sqrt{75}\] Do you agree with that?